### v1

# Li and Zhu

[…]
Li and Zhu
[…]
Human-centered Decision Making through Generative Curation
[…]
Balancing Optimality and Diversity:
Human-Centered Decision Making through
Generative Curation
[…]
The surge in data availability has inundated decision-makers with an overwhelming array of choices. While existing approaches focus on optimizing decisions based on quantifiable metrics, practical decision-making often requires balancing measurable quantitative criteria with unmeasurable qualitative factors embedded in the broader context. In such cases, algorithms can generate high-quality recommendations, but the final decision rests with the human, who must weigh both dimensions. We define the process of selecting the optimal set of algorithmic recommendations in this context as human-centered decision making. To address this challenge, we introduce a novel framework called generative curation, which optimizes the true desirability of decision options by integrating both quantitative and qualitative aspects. Our framework uses a Gaussian process to model unknown qualitative factors and derives a diversity metric that balances quantitative optimality with qualitative diversity. This trade-off enables the generation of a manageable subset of diverse, near-optimal actions that are robust to unknown qualitative preferences. To operationalize this framework, we propose two implementation approaches: a generative neural network architecture that produces a distribution $\pi$ to efficiently sample a diverse set of near-optimal actions, and a sequential optimization method to iteratively generates solutions that can be easily incorporated into complex optimization formulations. We validate our approach with extensive datasets, demonstrating its effectiveness in enhancing decision-making processes across a range of complex environments, with significant implications for policy and management.
[…]
human-centered decision making, generative curation, qualitative diversity, human-in-the-loop, algorithmic advice
1 Introduction
With the explosion of data availability, human decision-makers faces an ever-growing set of information and decision choices. This phenomenon has spurred significant interest in utilizing data-driven machine learning (ML) methods to assist in human decision-making, and there is a quickly expanding body of literature dedicated to developing methods capable of recommending the optimal decision across a variety of scenarios based on a particular objective (e.g. Bertsimas and Kallus 2016, Kube et al. 2019, Bryan and Ventresca 2020).
However, in real-world applications, these systems rarely operate independently. Instead, they are often embedded within a decision-making framework where the human decision-maker remains the ultimate authority. This human-centered approach is essential in many applications for a multitude of key reasons:
  1. 1.
    Legal and Ethical Accountability: In many complex decision-making contexts including clinical environments, governments, and courts, human decision-makers bear the ultimate responsibility for the outcomes of their decisions. Consequently, they must review and endorse any decision proposed by an algorithm to ensure accountability.
  2. 2.
    Addressing Incomplete or Inaccurate Objectives: The true desirability of an action is often multi-dimensional, encompassing qualitative and hidden factors such as personal preferences, ethical considerations, and political feasibility, which a single quantitative objective may fail to capture fully. Furthermore, quantitative objectives can be plagued by issues like bias, measurement errors, and selection bias. In such cases, human decision-makers are better equipped to contextualize these objectives, considering the broader implications and potential caveats.
  3. 3.
    Preference for Human Judgment: Research consistently shows that people generally prefer human decision-makers, even when humans are prone to similar or greater mistakes compared to algorithms (Dietvorst et al. 2015). This preference is supported in numerous experiments across domains from criminal sentencing decisions (Binns et al. 2018) to medical decision-making (Longoni et al. 2019). A recent survey also found that, despite the convenience of AI-assisted agents, people overwhelmingly preferred human decision-makers to address their concerns, even when it involved longer wait times (CallVU 2023).
In this paper, we focus on developing decision recommendations while explicitly considering the context that these algorithmic suggestions are subject to human arbitration. We define this category of decision-making algorithms as human-centered decision making. As illustrated in Table 1, this category contrasts with common optimal decision making algorithms that seeks the best decision excluding any human agency (e.g. Zhao et al. 2012, Moodie et al. 2012, Zhu et al. 2017, Luedtke and van der Laan 2016, Athey and Wager 2021). We instead formulate a collaborative model where algorithms provide potential solutions that humans then evaluate, ensuring that algorithmic insights augment, rather than replace, human expertise, particularly in complex, nuanced decision scenarios.
Table 1: Types of Decision Making Systems
Human has final decision agency?
No Yes
Algorithm gives decision input? No N/A Human Decision Making e.g. Doctors, Policymakers
Yes Optimal Decision Making e.g. optimal treatment rules Human-Centered Decision Making e.g. modeling to generate alternatives, generative curation
Current approaches in human-centered decision making often utilize ad-hoc constraints to heuristically generate a range of quantitatively near-optimal actions for decision-makers to choose from (Chang et al. 1982, Greistorfer et al. 2008). Alternatively, researchers try to elicit multiple dimensions of human preference and present different solutions along a Pareto frontier (Masin and Bukchin 2008, Lin et al. 2022, Wu et al. 2024).
We propose a different approach by recognizing that the ultimate goal of any recommendation system within a human-centered decision-making framework should be to facilitate decisions that are deemed most desirable by the human decision-maker themselves. Therefore, the goal of any set of $m$ recommendations $A_{1},\cdots,A_{m}$ should be to maximize the human desirability $\ell(\cdot)$ of the best recommendation
$\max_{A_{1},\cdots,A_{m}}\max_{i=1,\cdots,m}\ell(A_{i}).$ ($\mathcal{P}^{0}$)
If the human desirability $\ell(\cdot)$ is known, the problem directly degenerates to the setting of optimal decision making, where the goal is to find a single recommendation $A$ that maximizes $\ell(A)$.
However, human desirability of a decision is often not only affected by measurable, quantitative objectives, but also elusive qualitative factors from the broader problem context. We demonstrate this through several motivating examples:
Example 1.1 (Non-pharmaceutical Interventions for Pandemic Control)
During the COVID-19 pandemic, governments worldwide implemented a wide range of non-pharmaceutical interventions to manage the spread of the virus. Policymakers faced the challenge of balancing numerous quantitative factors, such as pandemic spread, control effectiveness, and fairness, with qualitative considerations like population acceptability, implementation feasibility, and political implications. Quantifying certain qualitative aspects is nearly impossible, and even for quantitative measures, policymakers often struggle to precisely determine their relative importance to formulate concrete objectives. Therefore, the implemented actions are often quantitatively suboptimal yet reflects a range of real-world desirability factors that are difficult to consider in model formulations. (Bertsimas et al. 2022)
Example 1.2 (Police Districting)
In police districting, departments typically organize their patrol forces by dividing a city’s geographical region into multiple patrol areas, known as zones. Quantitatively, the objective is to find an optimal districting plan that minimizes workload variance across zones (Larson 1974, Larson and Odoni 1981, Zhu et al. 2020, 2022, Xing et al. 2023). However, qualitative factors such as political and economic considerations and the shape of the districts are challenging to fully capture in a mathematically well-defined manner (Shirabe 2009). This complexity renders a comprehensive quantitative assessment of the underlying desirability of actions nearly impossible (Gardner et al. 2014). Figure 1(c) demonstrates a case where Atlanta police department rejected a quantitatively optimal plan due to ease of access to some highways and traffic constraints within the boundaries.
[…]
(a) Pre-2019 plan
(b) Adopted plan
(c) Rejected plan
An example showing the difficult-to-define qualitative desirability through Atlanta police redistricting. Gray lines represent the basic geographical units patrolled by the police, the color depth indicates the police workload of each unit. Black lines outline the districting plans, and dashed red lines highlight the changes made to the pre-2019 plan in (a). The map in (b) is the new plan adopted by the APD after 2019. The map in (c) is quantitatively optimal (the workload variance across zones is minimized) according to an optimization model but was ultimately rejected by the APD because it overlooked traffic constraints and inadvertently cut off access to some highways with its zone boundaries.
Therefore, human desirability of any decision includes both explicitly known quantitative objectives and also unknown qualitative desirability factors. When we incorporate both known and unknown factors into $\ell(\cdot)$, Problem $\mathcal{P}^{0}$ is distinct from the classic optimal decision making problem.
To solve Problem $\mathcal{P}^{0}$, instead of directly maximizing over the set of recommended decisions – which becomes computationally prohibitive, especially when $m$ is large – we employ a distributional approach where we aim to create the optimal generative distribution $\pi$ to sample recommendations. In other words, the generative distribution $\pi$ should maximize the human desirability of the best action within a set of sampled recommendations sampled from $\pi$:
$\max_{\pi}~{}\mathbb{E}\left[\max_{A_{1},\cdots,A_{m}\sim\pi}\ell(A_{i})\right].$ ($\mathcal{P}^{*}$)
Utilizing a distributional strategy grants decision-makers the flexibility to generate any number of viable solutions, circumventing the need to directly model the vast decision space of potential actions. Figure 2(b) provides a graphical illustration of this method, which we denote as generative curation.
Under generative curation, we can show that when desirability is influenced by arbitrary observable and unobservable characteristics, the objective function can be modeled as optimizing a quantitative objective with an added noise term, characterized by a stationary stochastic process. Specifically, we demonstrate that by modeling this stochastic process as Gaussian, the optimization problem becomes a trade-off between quantitative optimality and qualitative diversity, governed by a novel diversity metric derived analytically from the Gaussian process. Our exploration of this diversity metric uncovers that achieving diversity does not merely equate to generating a wide spread of solutions. In fact, even mild assumptions imposed on the Gaussian process can lead to highly non-trivial dynamics in how solution diversity is induced.
To operationalize this framework, we propose two implementation approaches: ($i$) We develop a generative neural network architecture to directly output $\pi$ that can be utilized to sample high-quality solutions in practical decision-making scenarios; ($ii$) We demonstrate that the framework can also be approximated through a sequential optimization approach, where high-quality solutions are iteratively generated. The latter method is particularly adaptable to complex optimization problems, such as routing or assignment tasks, where a distributional formulation of the problem is more difficult. We validate our approach using both synthetic and real-world datasets, demonstrating its effectiveness in enhancing decision-making processes in complex environments.
Formally, our contributions are as follows:
[…]
(a) Optimal decision-making
(b) Generative Curation
Comparison between classic optimal decision-making and the proposed framework. (a) In classic optimal decision-making settings, optimization models typically focus only on observable quantitative factors with the goal of identifying a single optimal solution. (b) Generative Curation offers multiple thoughtfully curated solutions to human decision-makers, enabling them to choose the most desirable one while actively involving them in the decision-making process.
  1. 1.
    Human-Centered Decision Making with Generative Curation: We formalize the concept of human-centered decision making, and considers generating a set of recommendations to maximize human desirability of the action, which could include both quantitative factors and unmeasurable qualitative factors. We introduce the framework of generative curation where we utilize a distributional approach to enable flexible generation of diverse and near-optimal solutions.
  2. 2.
    Novel Diversity Metric: We establish that if qualitative factors are assumed to follow a stationary Gaussian process, their contribution to the objective function can be formally expressed as a new diversity regularization term. We further illustrate that the diversity metric is not equivalent to generating a wide spread of solutions, and how varying assumptions about the Gaussian Process influence the behavior of this diversity metric.
  3. 3.
    Flexible Implementation: To operationalize our framework, we offer two methods: ($i$) a generative neural network architecture that generates a distribution $\pi$ to efficiently sample a diverse set of near-optimal actions and ($ii$) a sequential optimization approach that iteratively selects solutions, allowing the framework to be integrated into complex optimization problems without requiring fundamental changes to existing structures.
  4. 4.
    Empirical Validation: We validate our framework using extensive synthetic and real datasets, demonstrating its effectiveness and practicality in enhancing decision-making across a range of complex environments.
1.1 Related work
Our research aims to enhance human-centered decision-making by generating a small set of solutions that effectively consider both quantitative and qualitative factors to optimize for human desirability. This approach intersects with several strands of literature.
1.1.1 Modeling to Generate Alternatives
”Modeling to generate alternatives” (MGA) is a well-established approach in decision-making that offers policymakers a set of diverse, near-optimal solutions, providing multiple perspectives on a given problem (Brill Jr et al. 1982). Traditionally, MGA research has concentrated on sequential optimization methods (Chang et al. 1982, Greistorfer et al. 2008, Ingmar et al. 2020, Delarue et al. 2023) and integer programming techniques (Trapp and Konrad 2015, Danna et al. 2007, Ahanor et al. 2024) to generate varied yet feasible solutions. In parallel, the recommendation systems literature has investigated integrating diversity objectives into classical algorithms (Puthiya Parambath et al. 2016, Castells et al. 2021). More recently, the Boltzmann distribution has gained popularity for promoting diversity by sampling from probabilistic models (e.g. Mann et al. 2020). Despite the widespread use of predefined metrics such as pairwise distance, entropy, or coverage to quantify diversity, these measures often lack robust justification as proxies for optimal diversity, as Delarue et al. (2023) highlighted. Our research seeks to bridge this gap by proposing a novel diversity metric grounded in a principled modeling framework that captures both observable and unobservable elements of human desirability.
1.1.2 Multi-objective Optimization
Another significant area of related work involves multi-objective optimization, where different objectives are quantified and analyzed using weighted sums (Chankong and Haimes 2008) or by selecting a diverse set of Pareto-optimal solutions (Masin and Bukchin 2008, Lin et al. 2022). Our approach extends this literature by incorporating qualitative factors that are difficult to quantify. By integrating these qualitative aspects into our optimization framework, we offer a more comprehensive approach to generating decision recommendations that better reflect real-world complexities.
1.1.3 Human-AI Interactions
More broadly, our work relates to the growing literature on human-AI interactions. Many works have focused on understanding the impact of machine learning (ML) algorithms that provide advice to human decision-makers and explore optimal strategies to combine human and algorithmic decision-making (Grand-Clément and Pauphilet 2024, Orfanoudaki et al. 2022, Chen et al. 2023).
These works predominantly assume that the algorithmic input is a singular recommendation, whether as a specific policy or action, that the human incorporates into their final decision-making process. We diverge from this approach by providing a set of potential recommendations to ensure that the recommendations can balance quantitative optimality with qualitative diversity to take into account unknown human preferences.
2 Generative Curation for Human-Centered Decision Making
In this section, we present our main framework of generative curation designed to assist human-centered decision making by offering a set of decisions that are optimally desirable for both quantitative and unknown qualitative factors.
2.1 Problem Setup
Consider the scenario where a human decision-maker needs to make a decision (or action) $a\in\mathcal{A}\subseteq\mathbb{R}^{d}$ from a large, possibly infinite set of potential actions. The goal is to maximize the human desirability of the actions, $\ell(a)$. Typically, the decision-maker can only formulate a partial part of desirability that depend on some observable features of the actions (and/or the underlying population) $X_{1}(a),\cdots,X_{p}(a)$. There are other important qualitative considerations $V_{1}(a)\sim\mathcal{P}_{V_{1}},\cdots,V_{q}(a)\sim\mathcal{P}_{V_{q}}$ that are unobserved and cannot be easily elicited from data. Without loss of generality, we assume that the underlying true desirability $\ell$ is influenced arbitarily by both components as followed:
$\ell(a)\coloneqq\ell(X_{1}(a),\cdots,X_{p}(a),V_{1}(a),\cdots,V_{q}(a)).$ (1)
The desirability in Equation (1) can be also written as:
$\ell(a)=\mathbb{E}_{V_{1},\cdots,V_{q}}[\ell(a)]+\left(\ell(a)-\mathbb{E}_{V_{1},\cdots,V_{q}}[\ell(a)]\right)\coloneqq Y(a)+U(a),$
where $U$ is a zero-mean stochastic process. If $X_{1}(a),\dots,X_{p}(a)$ are based on observable features, then $Y(a)$ is a deterministic objective function that encodes the quantitative desirability of the action, whereas $U(a)$ retains the variation due to the unobserved qualitative desirability. Figure 3(c)(a) gives an illustrative example of both quantitative and qualitative desirability in an one-dimensional action space.
As detailed in Section 1, the goal of generative curation is to create a probability distribution over the action space $\pi(a):\mathcal{A}\to[0,1]$ that enables the decision-maker to maximize the human desirability of the best action among $m$ sampled actions. We reproduce Problem $\mathcal{P}^{*}$ below:
$\max_{\pi}~{}\mathbb{E}\left[\max_{A_{1},\cdots,A_{m}\sim\pi}~{}\ell(A_{i})\right],~{}\text{where}~{}\ell(A_{i})=Y(A_{i})+U(A_{i}).$ ($\mathcal{P}^{*}$)
The value of $m$ is typically limited by practical constraints. For instance, in the police districting problem (Zhu et al. 2020, 2022), the decision-makers were asked to present no more than five districting plans due to time constraints.
This formulation effectively integrates human agency with algorithmic decision-making. The quantitative algorithm generates a policy $\pi$ that samples $m$ actions, from which the human decision-maker selects the one that best aligns with their preferences. We note here that this approach does not assume the decision-maker has precise knowledge of their preference function $\ell(A_{i})$. There is considerable literature that supports individuals often struggle to articulate their own preferences accurately, yet they are significantly better at making comparative judgments between options (e.g., Fischhoff 1991, Tversky and Kahneman 1974, Slovic 1995). This comparative capability is sufficient for selecting the best option among $m$ actions (e.g., through bubble sort as detailed in Astrachan (2003)).
[…]
(a) Desirability
(b) $\epsilon$-optimal uniform policy
(c) $\epsilon$ vs regret
An illustrative example of the true desirability of a human decision-maker. (a) The quantitative desirability is represented by a Gaussian function and the underlying qualitative desirability is drawn from a GP with zero mean and the exponential kernel covariance function. Choosing quantitatively optimal action ($A=0.5$) could result in sub-optimal desirability. (b) The $\epsilon$-optimal uniform policy is adopted. Larger the value of $\epsilon$, more “diverse” the policy $\pi_{\epsilon}$. (c) The optimal desirability is attained at $\epsilon\approx 0.24$ when $m=20$.
2.2 A Tight Reformulation with Gaussian Processes
To solve Problem $\mathcal{P}^{*}$, it is necessary to impose certain assumptions on the structure of the stochastic process $U(A_{i})$. Specifically, we assume that this process can be modeled as a stationary Gaussian process (GP) with a mean of zero and a covariance function $k(a,a^{\prime})$: {assumption} The qualitative desirability $U(a)$ can be modeled by a stationary GP with a mean of zero and a covariance function $k(a,a^{\prime})$. This assumption is quite general as GPs are widely recognized in the literature for their flexibility and capacity to model a broad range of stochastic processes. By assuming stationarity, we are utilizing the principle of indifference to concretize the notion that since qualitative desirability is unknown, it therefore should not apriori depend on what the action is.
Utilizing Assumption 2.2, we can bound the objective in Problem $\mathcal{P}^{*}$ as followed:
Proposition 2.1
Problem $\mathcal{P}^{*}$ admits the following upper and lower bounds:
$\max_{\pi}~{}\mathbb{E}\left[\max_{A_{1},\cdots,A_{m}\sim\pi}~{}\ell(A_{i})\right]$ $\leq\max_{\pi}\left\{\mathbb{E}\left[\max_{A_{1},\dots,A_{m}\sim\pi}(Y(A_{1}),\cdots,Y(A_{m}))\right]+\sigma\sqrt{1-\rho[\pi]}E_{m}\right\},$ (2)
$\max_{\pi}~{}\mathbb{E}\left[\max_{A_{1},\cdots,A_{m}\sim\pi}~{}\ell(A_{i})\right]$ $\geq\max_{\pi}\left\{\mathbb{E}_{A\sim\pi}\left[Y(A)\right]+\sigma\sqrt{1-\rho[\pi]}E_{m}\right\},$ (3)
where:
$\sigma^{2}$ $=k(a,a),$
$\rho[\pi]$ $=\frac{1}{\sigma^{2}}\mathbb{E}[k(A_{i},A_{j})]=\frac{1}{\sigma^{2}}\int\int k(a_{i},a_{j})\pi(a_{i})\pi(a_{j})\mathrm{d}a_{i}\mathrm{d}a_{j},$
$E_{m}$ $=\int^{\infty}_{-\infty}x\frac{\mathrm{d}}{\mathrm{d}x}\Phi(x)^{m}\mathrm{d}x,$
and $\Phi(x)$ is the cumulative distribution function (CDF) of the standard normal distribution. Both bounds are tight when $Y(\cdot)$ is a constant function or $m=1$.
The proof is detailed in Appendix 6. By assuming that the qualitative desirability follows a general GP, we derive both lower and upper bounds for the original objective. In the case where $m=1$, the term $E_{m}$ becomes zero, reducing both bounds to the classical optimal policy problem that disregards diversity:
$\max_{\pi}\mathbb{E}_{\pi}[Y(A_{i})].$
This outcome confirms that when the goal is to recommend only a single solution to the decision-maker, the unobservable qualitative desirability can be effectively ignored.
However, when multiple solutions are to be recommended, the objective naturally evolves into a trade-off between quantitative optimality and a second term representing solution diversity. Specifically, $\rho[\pi]$ captures the expected correlation between the qualitative desirability components $U(A_{i})$ and $U(A_{j})$, serving as a measure of similarity or concentration among the generated solutions. When $\rho[\pi]=1$, the solutions are perfectly correlated in terms of their qualitative desirability, indicating a lack of diversity. The term $\sigma\sqrt{1-\rho[\pi]}$ in both bounds therefore incentivizes minimizing this correlation to promote diversity among the solutions.
From a managerial standpoint, this insight underscores the importance of diversity when presenting multiple candidate actions to decision-makers. If all options are too similar in their unobservable qualities, the decision-maker’s ability to identify the most desirable action is diminished. The framework naturally justifies the heuristic need for diversity in generated actions: by offering a broader range of qualitatively distinct options, the likelihood increases that the decision-maker will find an action that aligns with their complex and potentially unarticulated preferences.
To further illustrate this behavior, we consider a toy example where a uniform policy is applied to all actions that are $\epsilon$-optimal, as illustrated in Figure 3(c)(b). This setup provides precise control over both quantitative optimality and qualitative desirability by simply varying $\epsilon$. A higher $\epsilon$ increases the size of the policy’s support, introducing more diverse actions that may diverge from purely quantitative optimality. Figure 3(c)(c) visually illustrates this trade-off, showing a convex regret curve where increasing diversity leads to greater regret once $\epsilon$ exceeds 0.24.
Both bounds in Proposition 2.1 further depend on the parameters $\sigma$ and $E_{m}$. We analyze the significance of each of these parameters:
  1. 1.
    The parameter $\sigma$ represents the variance of the qualitative desirability component. Intuitively, $\sigma$ captures the expected magnitude of the impact that unobserved qualitative diversity has on overall desirability. As $\sigma\to\infty$, the unobserved qualitative factors become increasingly dominant relative to the quantitative factors. In practical applications, $\sigma$ is typically unknown and must be calibrated to reflect the preferences and behaviors of various stakeholders.
  2. 2.
    The term $E_{m}$ quantifies the expected maximum of the qualitative desirability across $m$ samples. As $m$ increases, extreme value theory suggests that $E_{m}\approx\sqrt{2\log m}$ (a detailed derivation is provided in Appendix 7). Thus, offering more solutions can also lead to an increase in consideration of the qualitative factors, as the decision-makers have more solutions to choose from. However, it is important to note that this effect is less pronounced than that of the $\sigma$ parameter. Although $\sigma$ and $E_{m}$ are exchangeable in the lower bound in Equation (3), a doubling of $\sigma$ would be roughly equivalent to a fourth-power increase in the number of solutions. This suggests that the significance of qualitative desirability for each decision far outweighs the importance of the number of solutions provided to the decision-maker.
In the following sections, for simplicity, we would concentrate on the conservative lower-bound objective in Equation (3), and explore various theoretical properties.
2.3 Theoretical Results
To analyze the behavior of our conservative objective function in Equation (3), we would first explore some basic properties. Note that since $\sigma$ and $m$ are exchangable parameters in the lower bound, equivalent statements hold as $m\to\infty$.
Proposition 2.2
Let $\pi^{*}(m,\sigma)$ be an optimal solution to Equation (3) and $\ell^{*}(m,\sigma)$ be the associated optimal value. Further define the following quantities:
$y^{*}(m,\sigma)$ $:=\mathbb{E}_{A\sim\pi^{*}(m,\sigma)}[Y(A)],$
$\rho^{*}(m,\sigma)$ $:=\rho[\pi^{*}(m,\sigma)],$
$u^{*}(m,\sigma)$ $:=\sigma\sqrt{1-\rho^{*}(m,\sigma)}E_{m}.$
Then, the following statements are true:
  1. 1.
    $\rho^{*}(m,\sigma)$ and $y^{*}(m,\sigma)$ are monotonically decreasing functions of $\sigma$.
  2. 2.
    Assume that $\pi^{*}(m,\sigma)$ is unique for every $\sigma$. Then, as $\sigma\to\infty$, $\pi^{*}(m,\sigma)\to\pi_{m}^{*}$. In particular, either $\pi^{*}_{m}$ is a degenerate distribution, or $\pi^{*}_{m}$ satisfies, for some constant $c$:
    $\int k(a,a^{\prime})\pi_{m}^{*}(a^{\prime})\mathrm{d}a^{\prime}=c\;\;\forall a.$
Proposition 2.2 confirms a few intuitive facts. First, as $\sigma$ increases, the importance of qualitative desirability increases, and therefore the expected quantitative desirability of the solutions and the covariance of the resulting solution distribution also decrease, increasing the qualitative desirability. This suggests that as $\sigma$ increase, there is a smooth tradeoff between the quantitative and qualitative desirability.
Furthermore, the second statement characterizes $\pi^{*}$ as $\sigma\to\infty$. In particular, it suggests that the optimal solution distribution either degenerates into a point distribution, or it satisfies the integral equation above. We can therefore characterize the asymptotic solution distribution for some commonly used kernels:
Corollary 2.3
  1. 1.
    Assume the space of actions $\mathcal{A}$ is a compact subset of $\mathbb{R}$, and $U(a)$ follows the white noise kernel $k(a,a^{\prime})=\kappa\mathbbm{1}\{a=a^{\prime}\}$. Then we have:
    $\pi^{*}_{m}(a)=\frac{1}{|\mathcal{A}|}.$
  2. 2.
    Assume the space of actions $\mathcal{A}$ is $[-q,q]$. If $U(a)$ follows the Gaussian kernel $k(a,a^{\prime})=\exp\left(-\frac{(a-a^{\prime})^{2}}{2h^{2}}\right)$, let:
    $\pi^{*}_{m}(a)=\frac{1}{2q}+\sum_{n=1}^{\infty}a_{n}\cos\left(\frac{n\pi a}{q}\right)$
    be the Fourier expansion of $\pi^{*}_{m}(a)$. Then the Fourier coefficients satisfy the following equation for all $\xi$ for some absolute constant $C$:
    $\frac{1}{2q}+\sum_{n=1}^{\infty}a_{n}2q\left(\operatorname{sinc}(2q\xi-\frac{n}{2q})+\operatorname{sinc}(2q\xi+\frac{n}{2q})\right)=\frac{C}{h}\delta(\xi)\exp(2h^{2}\pi^{2}\xi^{2}).$
The results in Corollary 2.3 formalizes the intuition based on different assumptions on the behavior for the unknown qualitative desirability:
  1. 1.
    The white noise kernel represents a scenario where the qualitative desirability of any random solution is entirely unpredictable. Here, the optimal asymptotic distribution is uniform, meaning that each solution is equally likely to be selected. This outcome aligns with the managerial insight that in highly uncertain environments, where qualitative factors are perceived as random, it may be prudent to diversify options equally across the action space. This strategy minimizes the risk of overlooking potentially optimal solutions by ensuring a broad exploration of the solution space.
  2. 2.
    The Gaussian kernel is commonly employed in GP and assumes that the covariance between solutions decreases exponentially as their distance increases. This assumption reflects the belief that the decision-maker is likely to have similar desirability for solutions that are close to each other – a concept supported by research in psychology that suggests proximity influences preference similarity (Shepard 1987, Nosofsky 1986). Under this kernel, the optimal asymptotic solution distribution is more complex. Numerically, they tend to form multiple clumps within the solution space, and in Appendix 10, we graph the optimal solutions for various values of $h$. This insight is crucial for managers, as it suggests that when decision-makers are likely to prefer similar options, the recommended solutions should be well-separated in multiple clumps across the action space.
Our results can be further connected to prior efforts aimed at inducing diversity in solution sets. Many works have focused on maximizing a pairwise distance metric to solve a variant of the following optimization problem (e.g., Greistorfer et al. 2008). The objective is to sample action $a_{i}$ that either heuristically or optimally maximize the pairwise distance (often in the $L_{2}$ metric), subject to the constraint that these actions should not deviate significantly from optimality:
$\max_{a_{i}}\sum_{i,j=1}^{n}\|a_{i}-a_{j}\|_{2}^{2},$ (4)
[…]
Considering the distributional version of this problem, we obtain:
$\max_{\pi}\;\;\mathbb{E}_{a_{i},a_{j}\sim\pi}[\|a_{i}-a_{j}\|_{2}^{2}],$ (5)
[…]
This leads to the following proposition:
Proposition 2.4
Let $\pi^{*}$ be the optimal solution to the problem in Equation (5). Assume $\mathcal{A}\subset\mathbb{R}$ and denote $a_{l}$ and $a_{u}$ as the smallest and largest values of $a$ that satisfy $Y(a_{i})\geq(1-\delta)Y(a^{*})$. Then, the optimal distribution is given by:
$\pi^{*}(a)=\frac{1}{2}\mathbbm{1}\{a=a_{l}\}+\frac{1}{2}\mathbbm{1}\{a=a_{u}\}.$
Proof 2.5
Proof: We have $\mathbb{E}_{a_{i},a_{j}\sim\pi}[\|a_{i}-a_{j}\|_{2}^{2}]=2\mathbb{V}_{\pi}(a)$, and the result follows immediately from the Bhatia–Davis inequality (Bhatia and Davis 2000).
This proposition reveals a counterintuitive outcome: If the goal is to maximize pairwise distance, the solution set tends to concentrate on just two specific solutions. This concentration, rather than promoting diversity, effectively reduces it by narrowing the range of recommended solutions. In particular, this result demonstrates the value of a principled approach towards generating diverse solutions for human decision-making.
2.4 Incorporating Human Preference Feedback
The framework presented above focuses on a common scenario where the algorithm generates a range of desirable actions for the human decision-maker to select from. We can further extend this framework to incorporate the feedback from human decision-makers, which allows the algorithm to refine its understanding of implicit preferences and enhance the accuracy of future recommendations in a multi-stage process.
For example, in a healthcare setting, the proposed framework could be used to present a subset of potential treatment plans to a physician. By analyzing which treatment plan the physician prefers over another, the algorithm can gain insights into considerations that may not be explicitly documented, such as patient-specific factors or the physician’s personal experience. This feedback loop enables the algorithm to refine its model of qualitative desirability, thereby improving the alignment of future recommendations with human expertise and intuition.
[…]
An illustrative example demonstrating how human preferences reshape the qualitative desirability. The solid line and shaded region indicate the mean and the 95% confidence interval of the qualitative desirability $U(A)$, respectively. In this example, we choose $k(a,a^{\prime})=\exp\left(-(a-a^{\prime})^{2}/2\right)$ and receive three human preferential decisions sequentially, i.e., $U(0.2)>U(0.5)$, $U(0.3)>U(0.7)$, and $U(0.1)>U(0.4)$.
To illustrate, consider two actions, $a_{1}$ and $a_{2}$, generated by policy $\pi$. For notational simplicity, we assume $Y(a)$ is a known constant and we only focus on the estimation of $U(a)$. A human decision-maker reviews these actions and expresses a preference. For instance, if $a_{1}$ is preferred over $a_{2}$, it means $U(a_{1})>U(a_{2})$. This feedback allows the qualitative desirability to be updated using its posterior distribution $U(a^{\prime})|U(a_{1})>U(a_{2})$ for any new test point $a^{\prime}$.
Given our setup under Assumption 2.2, this posterior distribution can be analytically derived and evaluated efficiently:
Proposition 2.6
Let $U_{1}=U(a_{1})$, $U_{2}=U(a_{2})$ represent the qualitative desirability of two proposed actions. Define $D\coloneqq U_{1}-U_{2}$ as the difference in qualitative desirability between these actions. Assuming that the human decision-maker prefers action 1 over action 2 ($D>0$), the qualitative desirability of any future action $U^{\prime}\mid D>0$ follows a conditional distribution: $\mathcal{N}\left(\mu_{U^{\prime}|D>0},\sigma^{2}_{U^{\prime}|D>0}\right),$ where
$\mu_{U^{\prime}|D>0}=$ $~{}\mu_{U^{\prime}}+\frac{\text{Cov}(U^{\prime},D)}{\sigma_{D}}\frac{\phi(\alpha_{D})}{1-\Phi(\alpha_{D})},$
$\sigma^{2}_{U^{\prime}|D>0}=$ $~{}\sigma_{U^{\prime}}^{2}-\frac{\text{Cov}(U^{\prime},D)^{2}}{\sigma_{D}^{2}}+\left(\frac{\text{Cov}(U^{\prime},D)}{\sigma_{D}^{2}}\right)^{2}\sigma_{D}^{2}\left(1-\frac{\phi(\alpha_{D})}{1-\Phi(\alpha_{D})}\left(\frac{\phi(\alpha_{D})}{1-\Phi(\alpha_{D})}+\alpha_{D}\right)\right),$
$\sigma_{D}^{2}=$ $~{}k(a_{1},a_{1})+k(a_{2},a_{2})-2k(a_{1},a_{2}),$
$\text{Cov}(U^{\prime},D)=$ $~{}k(a^{\prime},a_{1})-k(a^{\prime},a_{2}).$
Here $\alpha_{D}=\mu_{D}/\sigma_{D}$, $\phi(\cdot)$ is the standard normal probability density function, and $\Phi(\cdot)$ is the standard normal cumulative distribution function.
The proof is contained in Appendix 11. In Figure 2.4, we present an illustrative example demonstrating how the qualitative desirability $U(A)$ is updated when multiple human preferential decisions are received sequentially. With each new binary preference signal, the shape of the desirability function becomes clearer, and the uncertainty surrounding $U(A)$ diminishes.
This adaptive approach not only enhances the algorithm’s ability to align with human expertise but also ensures that the recommendations remain relevant and robust in dynamic, real-world environments.
3 Implementation of Generative Curation
This section presents two specific implementations of our proposed generative curation framework.
3.1 Deep Generative Approach
In recent years, there has been a growing interest in applying generative frameworks to complex optimization problems, driven by advancements in machine learning and artificial intelligence. Generative models, particularly generative adversarial networks (GANs) (Goodfellow et al. 2014), variational autoencoders (VAEs) (Kingma 2013), and their extensions, have shown significant potential in generating high-quality, diverse solutions across various domains. These models have been successfully applied in areas ranging from design optimization (Oh et al. 2019, Guo and Vogiatzis 2020), synthetic data generation (Xu et al. 2019), to black-box optimization (Krishnamoorthy et al. 2023), demonstrating their versatility and power in solving difficult optimization problems.
Input: Quantitative desirability $Y$; Covariance kernel $k$; Variance of qualitative desirability $\sigma$; Batch size $n$; Randomly initialized model parameter $\theta_{0}$; Variance of the noise $\sigma^{2}_{\text{NN}}$; Number of iterations $T$; Number of policy recommendations $m$; Learning rate $\alpha$;
Output: A set of generated actions $\{a_{1},\dots,a_{m}\}$, where $a_{i}=\phi(\epsilon_{\text{NN}};\theta_{T}),\epsilon_{\text{NN}}\sim\mathcal{N}(0,\sigma^{2}_{\text{NN}})$;
for $t=0,\dots,T$ do
       for $(i,j)$ in $\{(1,1),\dots,(1,m),\dots,(n,m)\}$ do
             Sample $\epsilon_{\text{NN}}\sim\mathcal{N}(0,\sigma^{2}_{\text{NN}})$;
             $a_{i,j}(\theta_{t})\leftarrow\phi(\epsilon_{\text{NN}};\theta_{t})$;
       end for
      $\theta_{t+1}\leftarrow\theta_{t}+\alpha\nabla_{\theta_{t}}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{1}{2m}\sum_{j=1}^{2m}Y(a_{i,j}(\theta_{t}))+\sigma\sqrt{1-\frac{1}{m}\left(\sum_{j=1}^{m}k(a_{i,2j}(\theta_{t}),a_{i,2j-1}(\theta_{t}))\right)}E_{m}\right)\right\}$;
end for
Algorithm 1 Generative Curation via Neural Network
Our proposed framework allows a wide range of generative models to directly parameterize the policy $\pi$, enabling their use in solving the optimization problem presented in Problem $\mathcal{P}^{*}$. By harnessing the capabilities of cutting-edge generative models, our approach can produce a rich distribution of actions that is difficult to capture with traditional parametric models. These generative models not only optimize the objective function but also maintain a desirable level of diversity, which is critical in many real-world applications. In the experiments, we focus on a simple yet flexible generative architecture using the reparameterization trick (Kingma 2013) for clarity and ease of implementation. Specifically, instead of sampling $a$ directly from the distribution $\pi$, we express $a$ as a deterministic function of the model parameter $\theta$ and an auxiliary noise variable $\epsilon_{\text{NN}}$ drawn from a standard normal distribution. Thus, the generative process can be written as:
$a=\phi(\epsilon_{\text{NN}};\theta),\quad\epsilon_{\text{NN}}\sim\mathcal{N}(0,\sigma^{2}_{\text{NN}}),$
where $\sigma^{2}_{\text{NN}}$ is the variance of the input noise, selected based on empirical convergence, and $\phi$ is a deep neural network parameterized by $\theta$ that transforms the input noise into samples drawn from the target distribution. To learn the model parameter $\theta$, we begin by generating actions through feeding random noise into the network, followed by optimizing the network parameters according to Equation (3) using gradient descent. The learning and generation procedures are summarized in Algorithm 1.
While we explore this particular model in detail, it is important to note that our framework is general enough to incorporate a broader range of generative models, including VAEs, GANs, Diffusion Models, and others. However, a thorough exploration of these additional models is beyond the scope of this paper.
Input: Quantitative desirability $Y$; Covariance kernel $k$; Variance of qualitative desirability $\sigma$; Number of kept samples $n$; Variance of the noise $\sigma^{2}_{\text{DIS}}$; Number of iterations $T$; Number of policy recommendations $m$;
Output: A sequence of generated actions $B_{T}$;
for $t=1,\dots,n$ do
       $a_{t}^{*}\leftarrow\arg\max_{a}Y(a)+\epsilon_{\text{DIS}}(a),\quad\epsilon_{\text{DIS}}(a)\sim\mathcal{N}(0,\sigma^{2}_{\text{DIS}})$;
end for
$B_{n}\leftarrow(a_{1}^{*},\cdots,a_{n}^{*})$;
for $t=n,\dots,T$ do
       $a_{t}^{*}\leftarrow\arg\max_{a}\left\{Y(a)+\sigma\sqrt{1-\frac{1}{n}\left(\sum_{j=1}^{n}k(a_{j}^{*},a)\right)}E_{m}+\epsilon_{\text{DIS}}(a)\right\},\quad\epsilon_{\text{DIS}}(a)\sim\mathcal{N}(0,\sigma^{2}_{\text{DIS}})$;
       $B_{t}\leftarrow(a_{t-n+1}^{*},\cdots,a_{t}^{*})$;
end for
Algorithm 2 Generative Curation via Diversified Iterative Search
3.2 Iterative Curation Approach
Despite the success of generative models in various fields, applying these techniques to integer optimization problems – such as routing, assignment, and scheduling – presents unique challenges. These problems are often characterized by discrete and combinatorical structures, where the solution space is non-continuous and highly constrained. Traditional methods in integer programming focus on identifying a single optimal solution, often using techniques such as branch and bound (Land and Doig 1960) or cutting planes (Gomory 1958). These methods are not inherently designed to generate a distribution of solutions, making the direct application of generative curation frameworks less straightforward.
To address this limitation, we propose an iterative approach that enables the generation of diverse solutions over multiple iterations. As illustrated by Figure 4(c), this approach allows the model to incrementally refine the solution set, ultimately converging to a distribution that can approximate the optimal solution distribution for Problem $\mathcal{P}^{*}$. Through each iteration, the generated solutions increasingly reflect the desired balance between diversity and optimality, leveraging the strengths of generative methods while addressing the unique constraints posed by complex optimization problems. Specifically, we consider an iterative approach as outlined in Algorithm 2. At every step, we optimize the sample estimate of the objective function in Equation (3). As shown in Figure 4(c), with well-chosen noise variance $\epsilon^{2}_{\text{DIS}}$ (similar to the approach used for selecting $\epsilon^{2}_{\text{NN}}$), this procedure can empirically converge to the same distribution as our optimal solution to the original problem.
4 Experiments
In this section, we evaluate the theoretical characteristics and practical effectiveness of our proposed framework using both synthetic and real-world examples. To achieve this, we test the two types of proposed generative policies:
  1. 1.
    We implement our generative model as indicated in Algorithm 1, and denote it as Neural Net Generative Curation (NN-GC). The model is parameterized by a simple three-layer and fully-connected neural network, where each of two hidden layers contains $1,000$ nodes. We set $n=64$, $T=500$, the dimension of the noise is $10$, and its variance is $\sigma^{2}_{\text{NN}}=0.1$.
  2. 2.
    We implement the iterative approach as detailed in Algorithm 2, and denote it as Diversified Iterative Search for Generative Curation (DIS-GC). The model is parameterized by a simple three-layer and fully-connected neural network, where each of two hidden layers contains $1,000$ nodes. We set $n=50$, $T=1,000$, $\sigma^{2}_{\text{DIS}}=2\times 10^{-2}$, and take last $m$ actions in the state $B_{T}$ as the output of the model. This approach aims to maximize the lower bound (3) of the original optimization problem, and is broadly compatible with classic optimization solvers.
We compare our methods’ performance against the following baselines:
  1. 1.
    A random strategy (Random), which ignores every quantitative metric and uniformly generates actions from the action space.
  2. 2.
    A Quantitative Optimizer (QO) that finds optimal actions, simply based on quantitative desirability $Y(\cdot)$ as the optimization objective. Specifically, stochastic gradient descent is employed as the solver for continuous settings, while simulated annealing is used for discrete settings.
  3. 3.
    A noisy version of QO that adds a small random noise to the quantitative objective to encourage exploration in the action space.
  4. 4.
    An Iterative Search algorithm (IS) that introduces the diversity to the solution sets by maximizing the pairwise distances between solutions based on Equation (4).
To ensure a fair comparison, we rerun each baseline $m$ times and select the action that gives the highest human desirability.
To measure the performance of the algorithms, we adopt total regret as the key evaluation metric in our synthetic experiments, defined as:
$R(\pi)=\ell(a^{*})-\max_{a_{1},\dots,a_{m}\sim\pi}\ell(a_{i})\geq 0,$
where $a^{*}$ is the most desirable action according to the true $\ell(\cdot)$.
To evaluate the robustness of the methods, we also examine the $95\%$ quantile upper bound (Up) and $5\%$ quantile lower bound (Low) of the regret. Furthermore, we measure empirical diversity using $\sqrt{1-\widehat{\rho}[\pi]}$, where $n$ random actions $\{a_{i}\}_{i\leq n}$ are generated from policy $\pi$, and $\widehat{\rho}[\pi]$ is computed as $1/\sigma^{2}\sum_{i=1}^{\lfloor{n/2}\rfloor}k(a_{2i},a_{2i-1})/\lfloor{n/2}\rfloor$.
4.1 Synthetic Results
Table 2: Performance on three synthetic data sets.
Methods $1$D Gaussian $2$D Ackley Knapsack
$\mathbb{E}R(\pi)$ Low Up $\sqrt{1-\widehat{\rho}[\pi]}$ $\mathbb{E}R(\pi)$ Low Up $\sqrt{1-\widehat{\rho}[\pi]}$ $\mathbb{E}R(\pi)$ Low Up $\sqrt{1-\widehat{\rho}[\pi]}$
NN-GC .005 .001 .012 .773 .573 .189 .972 .875
DIS-GC .002 .000 .002 .672 .695 .087 1.26 .849 .049 .000 .100 .936
QO .261 .034 .392 .085 1.70 1.04 2.51 .004 .696 .224 1.20 .000
QO+Noise .159 .003 .364 .189 1.44 .905 2.21 .295 .440 .187 1.17 .518
IS .110 .021 .451 .771 1.41 .991 2.13 .901 .424 .167 1.71 .781
Random .215 .005 .984 .946 2.21 1.39 3.08 .979 1.31 .498 2.13 .623
[…]
An example of the synthetic $2$D desirability function in a bounded action space. The transparent surface in red represents the true desirability $\ell$ and the surface in gray represents the quantitative desirability $Y$, which is an $2$D negative Ackley function. In this example, we set $\sigma=10$.
We construct three synthetic settings to quantitatively analyze the regret as well as diversity of all methods:
  1. 1.
    1D Gaussian: We choose a simple Gaussian function in a bounded 1D space $[0,1]$ as the quantitative desirability function $Y(\cdot)$; We set $\sigma=0.25$ and assume the qualitative desirabiity $U$ follows a zero-mean GP with its covariance function being an exponential kernel, i.e., $k(a,a^{\prime})=\exp\left(-(|a-a^{\prime}|^{2})/2h^{2}\right)$, where $h=1$.
  2. 2.
    2D Ackley: We select a negative Ackley function (Ackley 2012) in a bounded $2$D space $[-3,3]\times[-3,3]$ as the quantitative desirability $Y$ and utilize zero-mean exponential-kernel GP with $h=0.5$ for the qualitative desirability. Figure 4.1 illustrates an example of potential desirability functions with $\sigma=10$, where the quantitatively optimal action is $a^{*}=(0,0)$.
  3. 3.
    Knapsack: We apply the proposed framework to a discrete setting by solving the classic knapsack problem (Kellerer et al. 2004). This problem consists of a set of $d=10$ items, each characterized by a specific weight and value (randomly generated integers between 00 and $10$), and a knapsack with a defined capacity $20$. The objective is to determine the optimal combination of items that maximizes the total value without exceeding the knapsack’s capacity. We set $\sigma=10$ and model the qualitative desirability using a GP, where the Hamming distance is used to measure the “similarity” between two actions, and an exponential covariance kernel with $h=0.5$ is adopted. Each possible action is represented as a binary vector, where each element indicates whether a corresponding item is included ($1$) or excluded (00) from the knapsack. The total number of possible solutions is $2^{d}$, where $d$ is the number of items, reflecting the binary decision (included or excluded) for each item. More details on the experimental setup for the Knapsack problem can be found in Appendix 12.1. It is important to note that NN-GC is not readily applicable in this discrete setting because generating discrete samples with a generative model requires further modeling considerations that are beyond the scope of this paper.
Each experiment is repeated $50$ times for every method, and we calculate the average regret, along with the $5$% lower bound and $95$% upper bound, as well as the associated diversity of the suggested actions. For each trial, we sample one possible qualitative desirability function from $U$ and evaluate the regret for each policy with $m$ set to $20$. The synthetic results are summarized in Table 2.
[…]
(a) $1$D Gaussian
(b) $2$D Ackley
(c) Knapsack
Convergence of iterative curation in both regret and diversity. We implement iterative curation following Algorithm 2, with parameters set as $m=20$, $n=50$, $T=1,000$, and $\sigma^{2}_{\text{DIS}}=2\times 10^{-2}$. We denote the policy represented by the iterative approach at the $t$-th iteration as $\pi_{t}$ and assume the last $m$ actions in the state $B_{T}$ are “drawn” from $\pi_{T}$.
Overall, our proposed methods demonstrate significant improvements over the baseline approaches. Both NN-GC and DIS-GC yield notably lower expected regret while exhibiting reduced uncertainty in regret compared to the baselines. Our methods also achieve considerably higher diversity than QO and QO+Noise, though slightly lower than Random, underscoring the challenge of balancing diversity with optimality.
Additionally, NN-GC generally outperforms DIS-GC in higher-dimensional spaces, likely because DIS-GC’s iterative nature is more affected by the curse of dimensionality. Despite this, DIS-GC proves more versatile, as it can be effectively applied to problems like the Knapsack problem, where NN-GC struggles to perform. Figure 4(c) illustrates the convergence of DIS-GC for all three problems, demonstrating that both regret and diversity converge rapidly.
In Figure 5(d) and Figure 6(d), we analyze the performance of the proposed NN-GC method under varying $\sigma$ and $m$ values for the 2D Ackley function experiment. In line with our conclusion in Section 2.2 and 2.3, we observe that the $\sigma$ value is crucial in regulating the diversity of the optimal policy, with higher $\sigma$ values leading to more diverse generative policies. On the other hand, varying the number of generated actions $m$ is less effective in influencing the optimal policy compared to changes in $\sigma$. Notably, the optimal policy with larger $m$ tends to be slightly less diverse than with smaller $m$, as it can achieve the same level of qualitative desirability with less diversity.
[…]
(a) $\sigma=50$
(b) $\sigma=20$
(c) $\sigma=10$
(d) $\sigma=5$
Learned generative policies from the synthetic experiment with varying $\sigma$ in a $2$D space. The surface in blue indicates the PDF of learned policy, which is estimated by KDE using $10,000$ generated actions. The black dots represent $m=20$ actions suggested by the corresponding policy.
[…]
(a) $m=50$
(b) $m=100$
(c) $m=500$
(d) $m=1,000$
Learned generative policies from the synthetic experiment with varying $m$ in a $2$D space. We set $\sigma=10$ in this comparison. The PDF of learned policy (surface in blue) is estimated by KDE using $10,000$ generated actions. The black dots represent $m$ actions suggested by the corresponding policy.
4.2 Real Results
We also apply our proposed framework to a real-world dataset, focusing on the police redistricting problem in Atlanta, Georgia (Zhu et al. 2022). In large urban areas, police departments typically organize their patrol forces by dividing the city into multiple patrol zones (or precincts), with each zone further divided into smaller areas known as beats (or sectors) (Larson 1974). The configuration of these patrol zones influences both the demand for and capacity of police services within each zone and beat, as well as the travel times for patrol units. These factors together determine the police response times to emergency calls and crime incidents. Consequently, the design of patrol zones plays a critical role in the overall efficiency of police operations.
Since Atlanta’s police zones were last reconfigured in 2011, the city’s population has grown substantially, leading to increased police workload, particularly in high-growth areas such as North Atlanta (as shown in Figure 1(c)(a) and Figure 7(e)(a)). This population growth, coupled with recruitment challenges, has contributed to longer response times for high-priority 911 calls. To address these issues, we aim to reallocate $78$ geographical units within the Atlanta police system into six zones, with the goal of minimizing workload variance across the zones. The police redistricting problem can be formulated as a mixed-integer linear programming problem, where the decision variable is a $78\times 6$ binary matrix, with each entry representing the assignment of beat $i$ to zone $j$. For simplicity, the workload for each zone is calculated as the sum of workloads for all beats within that zone, and the workload of each beat is estimated using 911 calls-for-service data from 2013 to 2017. Because the real data lacks direct access to qualitative desirability, we model qualitative desirability $U$ as unknown to all models and use a GP with an exponential kernel, where $h=0.5$ and $\sigma_{U}=10^{2}$. Similar to the Knapsack problem, we use Hamming distance to measure the difference between two vectorized districting plans. The performance of all methods is assessed by measuring their best regret when $m=5$. Additionally, we extensively test our method with different variance values of $\sigma$ ranging from 00 to $10^{5}$, investigating how the choice of $\sigma$ affects regret in practice.
[…]
True regret of all methods on Atlanta police redistricting. We set $m=5$ and synthesize the underlying qualitative desirability using a GP with an exponential kernel, where $h=0.5$ and $\sigma_{U}=10$. We evaluate the true regret for all baseline methods as well as our DIS-GC with $\sigma=0,1,10,10^{2},10^{3}$.
[…]
(a) $16.77$ (Original)
(b) $7.99$
(c) $8.01$
(d) $8.72$
(e) $8.73$
The original and the new police redistricting plans generated by our proposed method. Black lines outline the suggested districting plans, and dashed red lines highlight the changes made to the pre-2019 plan. The color depth indicates the zone workload and the caption of each panel shows quantitative objective value (workload variance) of the corresponding plan. (a) is the original plan used before 2019, (b-e) are the plans randomly generated by our method. In particular, the plan in (c) resembles the adopted plan by APD shown in Figure 1(c) (b).
Figure 4.2 summarizes the performance of all methods. DIS-GC achieves the lowest regret and outperforms other baseline methods when $\sigma\approx 10^{2},10^{3}$. This demonstrates that our proposed framework can effectively recommend the most desirable actions when the variance $\sigma$ and kernel function $k$ are chosen appropriately (The selection of $\sigma$ and $k$ rely on human domain expertise in practice). We also present four random plans generated by DIS-GC with $\sigma=10$, alongside the original plan used by the APD before 2019, in Figure 7(e), all of which show significant reductions in zone workload variance. In addition to minimizing workload variance and ensuring zone contiguity, the police must consider several other practical factors that are not easily captured by mathematical models when selecting a final plan. For instance, the redesigned zones should be compact and aligned with the city’s transportation infrastructure, as patrol units typically travel within a zone. A narrow or irregularly shaped zone, or one divided by highways, can increase travel time and reduce patrol efficiency and emergency response effectiveness. Notably, plan (c) closely resembles the one implemented by APD in 2019 (shown in Figure 1(c) (b)), suggesting that this method could be applicable in real-world scenarios. Further details of the experimental setup and additional numerical results are provided in Appendix 12.2.
5 Discussion
In this paper, we presented a novel optimization framework termed Generative Curation to assist in human-centered decision making by optimizing the set of recommendations while explicitly considering human agency. Our approach integrates both measurable quantitative criteria and unmeasurable qualitative factors, addressing the limitations of traditional decision-making systems that primarily focus on optimizing based on quantifiable objectives. The framework leverages a GP to model the influence of unknown qualitative factors, facilitating a balance between quantitative optimality and qualitative diversity through a newly derived diversity metric. We proposed two implementation approaches: a generative architecture and a sequential optimization method, and we validate both using synthetic and real-world datasets to demonstrate their effectiveness in improving decision-making processes across various complex environments.
Despite the promising results demonstrated by our framework, we acknowledge that our assumption that qualitative desirability follows a stationary GP introduces assumptions about the nature of these factors that may not hold universally across different decision-making contexts. Stationarity might not hold true in cases where there is consistent, unknown bias for/against a certain decision region. The use of Gaussian distributions also limit the ability of the model to account for significant outlier deviations from the quantitative objective. We believe there is significant room for future work to better model unknown preferences.
We also note that within our current framework, we have assumed that the decision-maker would always select a decision within the recommended set. In some scenarios, decision-maker might only partially adopt one of the recommendations and make ad-hoc changes to the algorithmic recommendations. We believe there is significant more work to be done that can consider both the human agency of decision-making and the deviation of human decision maker from the set of recommendations.
In conclusion, while our generative curation framework offers a significant step in improving human-centered decision-making by integrating qualitative and quantitative factors, further work is needed to expand its applicability to a wider range of decision-making contexts.